sband wide/low (1.7 to 3.1 Ghz receiver)
The sband wide (sbw also called sband
low) receiver covers 1.8 to 3.1 ghz. It has a quadridge OMT
(native linear polarization) and a ying/kildal horn. The receiver
was installed on the telescope in 1999 (not sure of these dates). It
never worked very well because of the strong rfi at 1.940 Ghz
(cellular phones) and the 2.4-2.5 ghz pcs band. A filter bank
was built and installed that had filters at 1.7 -1.88,
2.04-2.4, and 2.6-3.1 ghz. With the filter bank installed, the
receiver started to work reliably (out of saturation!!).
- 23mar17: Tsys PolB jumps by 10K (more info)
- 13jan13: updated pointing offsets.
- 29sep08: installed
cal values. backdated to 18aug05.
- 10oct04: installed new cal values. backdated to 15aug04.
- 10sep04: new gain curve installed and back dated to be valid
- 14may04: cal diodes replaced.
- 14may04: horn moved down in focus direction to put it in
- 29apr04: horn moved x,y direction after feed survey (focus
motion not yet done).
- 1999-14may2004 The measured
temperature drifted from 1999 thru 14may2004. This was
caused by the cal diodes drifting. On 14may04 the diodes were
replaced and the measured system temperature stabilized.
calibration results: jun12-dec12. updated pointing offset.
26may08: cal values
measured using sky and absorber.
may04-sep04: sbw GAIN CURVES.
performance of data used for gain curves.
values measured using sky and absorber
values measured using sky and absorber
values measured using sky and absorber.
170817: Tsys PolB jumps by 10 K on 23mar17.
30may17: testing the correlated cal on sbw.
22sep15:retuned notch filters for 2130 rfi
11may15: notched filter to remove aws rfi at 2130
20jan11: rfi appears at 2380
for sbw dewar to cool down.
in the sbw receiver.
170817:Tsys polB > Tsys polA,
The Tsys for sbw polB has been higher than
Tsys PolA since 23mar17
plot of sbw tsysA,B difference
01jan17 thru 17aug17 (.ps) (.pdf)
- top: tsys (hcorcal) . black polA, red polB
- bottom: TsysA - tsysB
- Looking at the daily tsys output,
- 12mar->19mar No tsys measurements (hf heating campaign)
- 20mar17 : tsys, receiver is warm
- 22mar17: tsys polA 36K, polB47K, but some cal
configurations giving unrealistic values (tsys 16k)
- 23mar17: tsys PolA36k, polB 50K .. all cal configs
- I looked in the electronics receiver/instrumentation log
page to see why the receiver was warmed up, but the last
entry entry in the log was oct 2016 :)...
To see if the problem was the cal, or a real
problem with the system temperature, i looked at the calibration
runs for sbw during 2017. Unfortunately, calibration runs were
only done in june and july of 2017
- i usually check the dewar temperatures of a
receiver before scheduling a calibration run
- The dewar temp for sbw has been 999 for the entire year. Turns
out this was a problem with the temperature diode, and not a
problem with the receiver....
The gain and Tsys both use the cal value to convert to Kelvins.
For the SEFD (Tsys/Gain) the cal dependence cancels out.
- It the cal was a problem, Both Tsys and the gain for polB
would be high, but the SEFD for polB was be the same as
The plots show Tsys,gain, and
sefd for the jun,july, sbw calibration runs (.ps) (.pdf)
- Page 1: tsys and gain
- top frame: Tsys vs Za.
- + is polA, * is polB
- The colors are different frequencies:
- TsysA looks higher than tsysB
- 2nd Frame: TsysB/TsysA vs za
- The colors are different frequencies
- the purple line is a ratio of 1.
- Tsys Polb is 20 to 40% higher than polA
- The daily system temperature plots are taken at 2725MHz.
So the green + are the closest set.
- 3rd frame: Gain vs Za.
- + is polA, * is polB.
- The different colors are the different frequencies.
- bottom frame: GainB/GainA vs za
- the ratio looks to be greater than 1.
- The median values for the ratios:
- 2212: 1.02
- 2380: 1.09
- 2690: 1.04
- 2850: 1.08
- So the ratios are not close to the Tsys change (25%).
- Page 2: SEFD
- Top frame: SEFD polA,B
- + is polA, * is polB
- The colors are the different frequencies.
- Bottom frame: Ratio SEFD_B/SEFD_A vs za
- The median ratios are:
- 2212: 1.22
- 2380: 1.12
- 2690: 1.24
- 2850: 1.33
- Tsys polB for sbw increased from 40 to 50 K on 23mar17.
- This was after a receiver warm up, cool down.
- looking at the SEFDs:
- The ratio of the SEFD_B/SEFD_A is 1.24. This matches the 25%
increase in Tsys.
- So the problem is an increase in Tsys (not a problem with the
30may17 testing the correlated cal on sbw (top)
Joanna rankin reported trouble with the low
correlated cal on sbw. Her data set was:
- sbw high band
- low correlated cal, 25 hz winking cal
- 21apr14 (56768) and 04may14 (56781)
The data showed the winking cal was present,
but there was no correlation between polA and polB (as if
the 2 diode cal was used).
I checked the cima log files, and the low
correlated cal was requested...
On 30may17 i took some data using the mock spectrometer:
- sbw with high band filter installed
- rf centered at 2800 Mhz
- hcorcal with hardware winking cal selected.
- mock spectrometer:
- full stokes, 1024 channels, 1ms dumps,172 Mhz bw.
- A standard on was done with hardware winking cal, followed by
a cal on, off (not winking)
Processing the data:
- masgethwcal() was used to read the winking cal data.
- the cal on,offs were split into separate arrays (bwon,bwoff)
- the calon, caloff (10sec) files were input and the two recs
were stored in bcal
- the winking cal data (1ms resolution) was stored in
- masstokes() was used to use the calon,off to intensity and
phase calibrate the data.
- the cal values are used to intensity calibrate the data
- the polA,B phase differences from the system were fit and
removed using the atan of stokesU,stokesV ratio
The plots show the results of the
test (.ps) (.pdf):
- page 1. calon - calOff
- black is stokes I
- green is stokes Q
- red is stokes U
- blue is stokes V
- top: cal diff from the 10sec calon, caloff
- bottom: cal diff from the average of the 25hz winking cal
data (sampled at 1ms).
- Since the cal diode has the same path to polA and polB
(through couplers) it is the same as injecting the cal at an
angle of 45 degrees from the polA,polb pickups. So all of the
cal power should show up in stokes U...It does,..
- page 2:
- frequency average the calibrated 1ms sampled spectra..
giving total power time series.
- Top: stokes I
- Bottom: Stokes Q (green), stokes (U) red, stokes(V) blue
- You can see the 18K winking cal (9K polA,9KpolB).
- the green stokes q should be at zero, not -13K. This
is probably a result of bad cal values.
- the current cal values are from 2005. We remeasured sbw
back in 2016, but i never finished reducing the cal data
(too much rfi..)
- The correlated cals are working on the sbw receiver
- the constant , and the winking cal were both tested.
- I need to update the cal values for sbw.
22sep15: final notched filter for aws rfi (top)
We sent the sband notched filters back for
retuning. The first set did not fall off fast enough at the
right edge (2155Mhz).
On 22sep15 dana did a 2port network analyzer
measurement on the two filters we got back
plots show the 2 port measurements of the filters (.ps) (.pdf)
- Top frame: S21, S11 measurements of the two filters
- s21 (filter transmission):black filter a, red
- they are down 60db by 2110
- at 2155 they are still down by 55 db
- S11 (filter reflection). Black is filterA, red is filterB
11may15: notched filter for aws rfi (top)
We received the notched filter to remove the
AWS rfi around 2130 (more info
on the rfi)
The plots show the results of Dana's 2 port measurement
using the agilent network analyzer (.ps) (.pdf)
- S21 is the loss passsing through the filter
- S11 is the reflected power.
- S22, and S12 were similar
- I also plotted the phase (in deg) for S21
May04 to Sep04 fit GAIN
CURVES to calib data. (top...)
gain curve plot
Gain curves were fit to the sband wide
calibration gain data using 15may04 through 01sep04.
The start of this epoch was after the cals were replaced and the
horn had been lowered put it in focus. The plots show the
gain data (black) and the fits (red) for 2212, 2380, 2690, and
2850 Mhz. These gain equations were installed on 10sep04 and back
data to be valid starting on 15may04.
The fit used the za and the azimuth terms. We
probably could have use a little better az,za coverage. I included
the az terms (instead of just a za fit) because it looked like the
1az term had a significant amplitude.
- Fig 1 shows the az,za distribution for the data. The fit used
a linear fit in za up to za=14. Above 14 degrees terms in
(za-14)^2 and (za-14)^3 were included. The fit also
included 1az, 2az, and 3az sin, cos terms.
- Fig 2 plots the gain data and the fit to za. The fit
equation is printed with the sigma for the fit (in K/Jy).
- Fig 3 plots the fit residuals (data-fit) vs za.
The 2690 freq fit had a lot more scatter than the
other 3 frequencies. If you look at the system performance data for
this data set you'll see that the gain and cals were jumping around
but the sefd was solid. The problem is most likely being caused by
rfi in the cal measurement.
The routine gainget() or corhgainget() will
now return the sbw gain for data after 15may04 from these
equations. The coefficients can be found in the ascii file
data/gain.datR7 (this is provided in the AO idl
distribution for correlator routines). You can also find a copy of
it at AO in /pkg/rsi/local/libao/phil/data/gain.datR7.
Be careful using these routines for data before 15may04. There is
not gain curve and the routine should return and
status reflecting that...
may04 to sep04 :
System performance of data used to compute gain curves. (top...)
Heiles calibration scans done from
15may04 (new cals, horn lowered) thru 01sep04 were used to measure
system performance. This data was then used to compute the gain
curves for use on data taken after 15may04.
The first set of plots show
the system performance with all frequencies overplotted. The
sources are identified
by symbol and the frequencies by color.
The second set of plots has the
plotted separately for each frequency. The colors and symbols
are used to differentiate the sources. The figures are:
- Fig 1 shows the distributions on the dish of the measurements.
- Fig 2 has the Gain in Kelvins/Jansky. This relies on the cals
and the source flux. The next plot is Tsys vs za in
Kelvins followed by the SEFD (System Equivalent Flux
density) in Janskies / Tsys. At the bottom is the average beam
width in arc seconds.
- Fig 3 plots the coma parameter, first sidelobe height below
the peak, the main beam efficiency, followed by the main beam +
1st sidelobe beam efficiency.
- Fig 4 has the pointing errors in az,za.
The data for 2690 is jumping around a lot more than the other 3
frequencies. The gain and tsys jump but the sefd is relatively
stable. The cal measurement is probably getting rfi in the cal on or
the cal off causing tsys,gain to jump around.
- Fig 1 2212Mhz Gain,Tsys . 260 points
- Fig 2 2212 Mhz sidelobes,beam efficiencies
- Fig 3 2380 Mhz Gain,Tsys 260 points.
- Fig 4 2380 Mhz sidelobes,beam efficiencies.
- Fig 5 2690 Mhz Gain,Tsys 151 points.
- Fig 6 2690 Mhz sidelobes,beam efficiencies
- Fig 7 2850 Mhz Gain,Tsys 259 points.
- Fig 8 2850 Mhz sidelobes,beam efficiencies